OpenStudy (anonymous):

f(x)=7x+9 find f(6c-4) a. 35c + 9 c. 42c – 19 b. 42c + 9 d. 42c + 19

OpenStudy (hyper):

Answer: c or 42c -19 Solution: subbing x = 6c -4 into f(x), the expression become 7(6c -4) + 9 = 42c -28 + 9 = 42c -19